Use the arithmetic sequence of numbers 1, 3, 5, 7, 9,…to find the following:
a) What is d, the difference between any two consecutive terms?

Answer by jim_thompson5910(21685)

d=nth_term-previous_term
d=9-7=7-5=5-3=3-1=2

So the difference d is 2